# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

### v1

@@ Abstract
Distributed machine learning in high end-to-end latency and low, varying bandwidth network environments undergoes severe throughput degradation. Due to its low communication requirements, distributed SGD (D-SGD) remains the mainstream optimizer in such challenging networks, but it still suffers from significant throughput reduction. To mitigate these limitations, existing approaches typically employ gradient compression and delayed aggregation to alleviate low bandwidth and high latency, respectively. To address both challenges simultaneously, these strategies are often combined, introducing a complex three-way trade-off among compression ratio, staleness (delayed synchronization steps), and model convergence rate. To achieve the balance under varying bandwidth conditions, an adaptive policy is required to dynamically adjust these parameters. Unfortunately, existing works rely on static heuristic strategies due to the lack of theoretical guidance, which prevents them from achieving this goal. This study fills in this theoretical gap by introducing a new theoretical tool, decomposing the joint optimization problem into a traditional convergence rate analysis with multiple analyzable noise terms. We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression on training performance, filling a critical gap in understanding how compressed and delayed gradients affect training. Furthermore, by integrating the convergence rate with a network-aware time minimization condition, we propose DeCo-SGD, which dynamically adjusts the compression ratio and staleness based on the real-time network condition and training task. DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and static strategy in high-latency and low, varying bandwidth networks, respectively.
@@ 1 Introduction
Recent breakthroughs in large language models (LLMs) (e.g., GPT-4 gpt-4 , DeepSeek-v3 deepseek ) have been accompanied by an explosive growth in model parameters, leading to unprecedented communication demands for distributed training scalalaws2 ; ScalaLaws1 . State-of-the-art (SOTA) distributed training strategies flashattention ; megatron-lm2 are predominantly optimized for high-speed interconnects like PCIe 5.0 and NVLink nvidia , where bandwidth exceeds $10$ Gbps and latency remains under $10$ ms. Although these assumptions are reasonable in well-resourced data centers, they often fail to hold in dynamic network environments with high latency and low bandwidth jaghouar2024opendiloco . Unfortunately, such unfavorable network conditions are common in distributed training over wide area networks (WAN), such as federated learning dagctmc and cross-geographic distributed learning hsieh2017gaia . WAN-based distributed training lowers the barrier for LLM pre-training, empowering individual researchers and small labs while reducing the reliance on costly, specialized networking SIGCOMM24StellaTrain .
In these challenging dynamic network settings, distributed SGD (D-SGD) and its variants, such as FedAVG fedavg and D-QSGD cui2021slashing , remain the dominant optimization methods due to their communication-friendly nature. Despite its low communication requirements, D-SGD still suffers from significant throughput degradation under high-latency and low-bandwidth conditions. As illustrated in Fig. 1, when the bandwidth drops below $2$ Gbps and the latency exceeds $200$ ms, the throughput of D-SGD in training GPT-2 with four nodes reduces to $50\%$ of maximum achievable throughput [^1] The maximum achievable throughput is determined by the compute capacity of the machines, without network constraints.. {wrapfigure}t0.4 [Uncaptioned image]
[…]
@@
Gradient compression and delayed aggregation are common strategies to improve the throughput of D-SGD in a lossy manner, but joint usage introduces a complex three-way trade-off. Gradient compression [^2] In this work, the gradient compression algorithm carries vanilla error-feedback (EF) stich2019error by default. EF is a mechanism that collects and reuses the errors from the gradient compression to retain the model accuracy. Other EF strategies, such as EF21 EF21 or EControl Stich2023EControl are out of this paper’s scope. reduces communication overhead by compressing gradient information before transmission, to mitigate the low bandwidth problem dgc . Delayed aggregation DGA applies updates using gradients computed several iterations earlier to alleviate the high latency problem. To overcome the problem of low-bandwidth and high-latency at the same time in WANs, we need to use these two strategies jointly, introducing intricate trade-offs in two key dimensions: (1) Iteration-to-accuracy: The impact of compressing delayed gradients on convergence remains uncertain, as it is unclear to what extent aggregating delayed compressed gradients will affect model accuracy, due to the lack of theoretical analysis in prior work. (2) Time-to-iteration: The effect of staleness and compression ratio on communication-computation overlap is not fully understood. Specifically, the existence of a threshold compression ratio, below which further compression no longer improves per-iteration time, remains unexplored. This threshold varies dynamically with the bandwidth and staleness.
@@
To achieve optimal time-to-accuracy, we analyze the above two dimensions separately. First, in iteration-to-accuracy, we propose a new analytical tool called the Nested Virtual Sequence. This method builds on the original virtual sequence by constructing multiple virtual queues, breaking down the complex noise term into multiple analyzable noise terms. With this tool, we are the first to derive the convergence rate of distributed SGD with delayed aggregation and gradient compression (DD-EF-SGD). This convergence rate establishes the iteration-to-accuracy relationship and reveals that the value of staleness exponentially worsens the negative effect of the gradient compression on the model convergence. Second, we use mathematical induction to model the relationship between the end-to-end time and the number of iterations in DD-EF-SGD. In this way, we build the time-to-accuracy relationship and propose an adaptive algorithm (shown in Fig. 3) that can select the optimal compression ratio and staleness in the challenging dynamic network conditions, thus achieving the balanced three-way trade-off during the whole training.
In a word, the contributions of our work are summarized as below:
@@
  • We propose the Nested Virtual Sequence, a theoretical tool to analyze complex distributed SGD variants, by reducing it to a standard D-SGD process and several analytical noise terms. We are the first to derive the convergence rate of DD-EF-SGD with the tool, revealing that the delayed aggregation exponentially worsens the negative effect of the compression on the training process.
@@
  • We mathematically establish an approximate relationship between end-to-end time and accuracy. Based on this, we propose DeCo-SGD, a joint optimization algorithm that adaptively calculates the optimal compression ratio and staleness according to the training environments.
@@
  • We apply DeCo-SGD  in challenging networks environments (average bandwidth < $1$Gbps, latency $\geq$ $100$ms) across diverse model architectures including CNN, GPT and ViT. Experimental results validate the great three-way trade-offs of our design, achieving up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and the SOTA static strategy, respectively.
@@ 2 Preliminaries
We specifically concentrate on distributed machine learning (DML) in data-parallel mode using the DD-EF-SGD as the optimizer.
Figure 2: The running timelines for D-SGD and D-SGD variants. Both D-SGD and D-EF-SGD are serial processes, with D-EF-SGD achieving a reduced transmission time. Similarly, DD-SGD and DD-EF-SGD operate serially; while DD-SGD maintains the same communication time as D-SGD, DD-EF-SGD reduces transmission time.
@@ 2.1 The Optimization Problem of DML
In the data-parallel paradigm, the training dataset is partitioned across $n$ worker nodes, each holding a subset of the data. The optimization problem of DML aims to minimize a global loss function $f(\cdot)$ based on the model parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $f(\cdot)$ is defined as a weighted sum of local loss functions of the $i$-th worker $f_{i}(\cdot)$:
@@ 2.2.1 D-SGD
At $t$-th iteration, the $i$-th worker computes gradients $\mathbf{g}_{t}^{i}$ based on its local data and model parameters at $t$-th iteration $\mathbf{x}_{t}$. $\gamma$ is the stepsize and these local gradients are aggregated to update the global model, following $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. As shown in Fig. 2, D-SGD is a serial computation process. In this process, computation and communication take place sequentially in each iteration, and communication includes end-to-end latency and transmission time.
2.2.2 D-SGD with Gradient Compression and EF (D-EF-SGD)
D-EF-SGD stich2020communication enhances D-SGD with the sparsification gradient compression [^3] In this work, the sparsification compressor in D-EF-SGD is Top-$k$ by default, a well-performed compressor in many hybrid gradient compressors stc ; CocktailSGD . by incorporating EF mechanisms to compensate for sparsification compression errors. Sparsification compression is $\mathbf{e}_{t}^{i}$ represents the error term of the $i$-th worker at the $t$-th iteration, storing the discrepancy between the compressed gradient and the true gradient. We denote $\mathbf{C}_{\delta}(\cdot)$ as the compression operator with the compression ratio $\delta$, $\Delta_{t}^{i}$ as the compressed update at $t$-th iteration in $i$-th worker, and the update rules for D-EF-SGD are as follows:
$\begin{gathered}\Delta_{t}^{i}=\mathbf{C}_{\delta}(\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}),\quad\quad\quad\mathbf{e}_{t+1}^{i}=\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}-\Delta_{t}^{i},\quad\quad\quad\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\Delta_{t}^{i}.\end{gathered}$
As shown in Fig. 2, D-EF-SGD retains a serial process, but reduces transmission time compared to D-SGD.
2.2.3 D-SGD with Delay Aggregation (DD-SGD)
@@
DD-SGD DGA addresses communication delays by accounting for the delay staleness $\tau$ between the computation of gradients and their application to the model update stich2019error . $\mathbf{g}_{t-\tau}^{i}$ represents the gradient computed by worker $i$ at the $(t-\tau)$-th iteration and the update rule shows as $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}.$ The delay aggregation does not change the communication time for each iteration. As shown in Fig. 2, DD-SGD parallelizes computation and communication by performing the $t$-th computation at the $t$-th iteration and sending $t$-th gradient, but the aggregation applies the gradient computed at $(t-\tau)$-th iteration. When combined with gradient compression (DD-EF-SGD), the transmission time is further reduced, as shown in the last part of Fig. 2.
@@ 2.3 Virtual Sequence
Virtual Sequence stich2018sparsified ; mania2017perturbed is an analytical tool that simplifies the convergence analysis of D-SGD variants like D-EF-SGD. It defines an auxiliary sequence, $\tilde{\mathbf{x}}_{t}$, which evolves according to an idealized update rule and serves as a reference for the actual iterates, $\mathbf{x}_{t}$, accounting for noise introduced by compression and delay. We define the noise term as $B_{t}$, then we have $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$. We generalize updates as $\mathbf{v}_{t}$, (i.e., $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\mathbf{v}_{t}$) and $B_{t}$ can be computed in $B_{t+1}=B_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\mathbf{v}_{t}$. Then we have $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\textbf{g}_{t}^{i}$. This can help theoretically analyze D-SGD variants, but fails in complex variants like DD-EF-SGD, because the noise term decoupled from DD-EF-SGD cannot be analyzed. A new method is needed to analyze these complex D-SGD variants.

### v2

# Taming Latency and Bandwidth: A Theoretical Framework and Adaptive Algorithm for Communication-Constrained Training

@@ Abstract
Regional energy caps limit the growth of any single data center used for large-scale model training. This single-center training paradigm works when model size remains manageable, but exponential growth in the model size and computational demand challenges it. A natural alternative is to distribute training across multiple data centers over wide-area networks. This pools distributed resources, but suffers from high latency and low, time-varying bandwidth, sharply reducing throughout. Employing jointly gradient compression and delayed aggregation can alleviate communication problems, but introduces a complex three-way trade-off among compression ratio, staleness (delayed synchronization steps), and convergence rate. Existing work lacks theoretical guidance and can only propose fixed strategies, insensitive to computation and communication conditions. We address this with a new theoretical tool, decomposing the joint optimization problem into a traditional process plus multiple analyzable noise terms. Our analysis yields the first convergence rate for this setting and shows that increasing staleness exponentially amplifies the detrimental effect of compression. Leveraging these insights, we propose DeCo-SGD, which dynamically selects the compression ratio and staleness based on the real-time communication and computation conditions. DeCo-SGD achieves up to $5.07\times$ and $1.37\times$ speed-ups over distributed SGD and static strategy in high-latency and low, varying bandwidth networks, respectively.
@@ 1 Introduction
The rapid progress of large language models (LLMs) (e.g., GPT-4 gpt-4, DeepSeek-v3 deepseek) has been accompanied by an explosive growth in model parameters. This expansion drives a steep rise in the computational resources required for training scalalaws2; ScalaLaws1. Training in a single AI computing center, where thousands of GPUs are tightly coupled via high-speed interconnects (e.g., PCIe and NVLink nvidia), is increasingly constrained by regional energy capacity. This bottleneck is already evident in major hubs; for instance, data centers in Virginia are projected to consume up to $51\%$ of the state’s electricity by 2030 GEI-Datacenter-25, while utilities in Ohio have imposed moratoriums on new connections due to transmission limits Williams2024. As model sizes continue to grow exponentially, the paradigm of training in a hyperscale compute center is approaching its scalability limits. This observation motivates a new phase of large-scale training that extends beyond a single compute center or energy domain. Distributed training over wide-area networks (WANs) offers a practical and sustainable way to overcome regional power limits and coordinate multiple compute centers globally.
Recently, some pioneering projects have already demonstrated the feasibility of training over WANs as shown in Fig. 2. Intellect-1 collaboratively trained a 1010B-parameter model on up to $112$ H100100s distributed across five countries and three continents jaghouar2024opendiloco. Similarly, DeepLink connected datacenters across two cities in China, more than 1,5001{,}500 km apart, to train a model exceeding 100100B parameters, highlighting the potential of multi-region training over WANs to help alleviate regional power constraints and begin to harness geographically distributed compute at scale deeplink.
Distributed training over WANs is fundamentally communication-constrained, facing high latency and limited bandwidth. In this setting, the standard data-parallel paradigm requires workers to perform local computation and then exchange full gradients with the server. Then the server aggregates the gradients, updates the model, and broadcasts the new parameters to workers. Among algorithms employing this paradigm, Distributed Stochastic Gradient Descent (D-SGD) and its variants are the most widely adopted, thanks to their clean iterative formulation and practical scalability dagctmc; SIGCOMM24StellaTrain. However, under WAN conditions, the need to transmit full gradients leads to severe degradation in end-to-end training throughput (i.e., the number of samples processed per second) for these methods dong25emnlpsurvey.
Figure 1: Pioneering LLM training over WANs. DeepLink connects compute clusters in Shanghai and Ji’nan over 1,5001{,}500 km, with an estimated throughput efficiency of $35\%$. Intellect-1 trains across continents, where the corresponding throughput efficiency drops to approximately $15\%$.
Figure 2: The heatmap of throughput efficiency ($\%$) for D-SGD with four nodes training GPT-2, under different latency and bandwidth conditions. Each node is equipped with one A4040 GPU. The throughput efficiency at (x, y) is defined as the throughput at (x, y) divided by the maximum achievable throughput of the machines.
To quantify the impact of network conditions, we conduct some controlled experiments training GPT-2 under varying bandwidth and latency configurations. Fig. 2 shows the resulting throughput efficiency, defined as end-to-end training throughput normalized by the maximum achievable throughput [^1] This is determined by the compute capacity of the machines, without network constraints.. By mapping real-world parameters to this heatmap, we observe severe degradation: DeepLink (roughly 1,0001{,}000 Mb/s bandwidth and $100$ ms latency amazon) falls into a region with only about $35\%$ of the ideal throughput. Similarly, Intellect-1 (roughly $100$ Mb/s bandwidth and $500$ ms latency amazon) lies in a region with throughput efficiency around $15\%$. Taken together, this inefficiency underscores that overcoming the communication bottleneck is not merely an optimization, but a prerequisite for unlocking true scale of training across regions.
@@
Gradient compression and delayed aggregation are common strategies to improve the throughput of D-SGD in a lossy manner, but joint usage introduces a complex three-way trade-off. Gradient compression reduces communication overhead by compressing gradient information before transmission, to mitigate the low bandwidth problem yan2022acJsac. Delayed aggregation Chuanwu2022SAPipe; DGA applies updates using gradients computed several iterations earlier to alleviate the high latency problem. While combining these strategies offers a promising path to overcome the problem of low-bandwidth and high-latency at the same time, it introduces intricate trade-offs in two key dimensions: (1) Iteration-to-accuracy: The impact of compressing delayed gradients on convergence remains uncertain. It is unclear to what extent aggregating delayed compressed gradients will affect model accuracy, due to the lack of theoretical analysis in prior work. (2) Time-to-iteration: The effect of staleness and compression ratio on communication-computation overlap is not fully understood. Specifically, the existence of a threshold compression ratio, below which further compression no longer improves per-iteration time, remains unexplored. This threshold varies dynamically with the bandwidth and staleness.
@@
To minimize the wall-clock time required to reach the target accuracy, we analyze the above two dimensions separately. First, in iteration-to-accuracy, we propose a new analytical tool called the Nested Virtual Sequence. This method builds on the original virtual sequence by constructing multiple virtual queues, breaking down the complex noise term into multiple analyzable noise terms. With this tool, we are the first to derive the convergence rate of distributed SGD with delayed aggregation and gradient compression (DD-EF-SGD). This convergence rate establishes the iteration-to-accuracy relationship and reveals that increasing staleness exponentially worsens the negative effect of the gradient compression on the model convergence. Second, we use mathematical induction to model the relationship between the end-to-end time and the number of iterations in DD-EF-SGD. In this way, we build the time-to-accuracy relationship and propose an adaptive algorithm that can select the optimal compression ratio and staleness in the challenging dynamic network conditions, thus achieving the balanced three-way trade-off during the whole training.
In summary, the contributions of our work are as follows:
@@
  • We propose the Nested Virtual Sequence, a theoretical tool to analyze complex distributed SGD variants, by reducing it to a standard D-SGD process and several analytical noise terms. We are the first to derive the convergence rate of DD-EF-SGD with the tool, revealing that delayed aggregation exponentially worsens the negative effect of the compression on the training process.
@@
  • We mathematically establish an approximate relationship between end-to-end time and accuracy. Based on this, we propose DeCo-SGD, a joint optimization algorithm that adaptively determines the optimal compression ratio and staleness under varying training environments.
@@
  • We apply DeCo-SGD  in challenging network environments (average bandwidth $\leq$ 11 Gbps, latency $\geq$ $100$ ms). We conduct extensive experiments that cover diverse model architectures (ResNet, GPT, ViT), various network conditions, and system scalability. Experimental results demonstrate that our design achieves up to $5.07\times$ and $1.37\times$ speed-ups over D-SGD and the SOTA static strategy, respectively.
@@ 2 Preliminaries
In this work, we focus on distributed machine learning (DML) in data-parallel mode using the DD-EF-SGD as the optimizer.
@@ 2.1 The Optimization Problem of DML
In the data-parallel paradigm, the training dataset is partitioned across $n$ worker nodes, each holding a subset of the data. The optimization problem of DML aims to minimize a global loss function $f(\cdot)$ based on the model parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $f(\cdot)$ is defined as the average sum of local loss functions of the $i$-th worker $f_{i}(\cdot)$:
@@ 2.2.1 D-SGD
At the $t$-th iteration, the $i$-th worker computes gradients $\mathbf{g}_{t}^{i}$ based on its local data and model parameters at $t$-th iteration $\mathbf{x}_{t}$. $\gamma$ is the stepsize and these local gradients are aggregated to update the global model, following $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. As shown in Fig. 3, standard D-SGD follows a strict sequential execution. The computation phase starts until the communication phase is fully completed, and communication includes end-to-end latency and transmission time.
2.2.2 D-SGD with Gradient Compression
Gradient compression is a typical communication optimization method to reduce the communication volumes during the training process by compressing transferred gradients cui2021slashing. Existing strategies generally fall into three categories: sparsification, quantization, and low-rank compression grace. Quantization compression lQuant2025jsac reduces the communication volume by decreasing the number of bits of gradients. However, it struggles to achieve aggressive compression ratios (such as $1\%$ or less) and lacks fine-grained adaptivity (like changing ratios from $12\%$ to $10\%$ or $4$-bit into $4.5$-bit quantization) due to its discrete value selection. Low-rank compression vogels2019powersgd is tailored for the all-reduce communication primitive and well suited for training within a single data center. However, it incurs higher compression costs compared to other compression strategies and cannot leverage its advantages in decentralized training, which prefers peer-to-peer communication under delayed aggregation strategy DGA. Sparsification compression transmits dagctmc only parts of gradients, offering two advantages for decentralized distributed training: (1) It enables aggressive compression of $1\%$, with the help of the error feedback (EF) mechanism [^2] In this work, the gradient compression algorithm carries vanilla EF stich2019error by default. EF is a mechanism that collects and reuses the errors from the gradient compression to retain the model accuracy. Other EF strategies, such as EF21 EF21 or EControl Stich2023EControl are out of this paper’s scope.; (2) Compression ratios are continuously selectable, facilitating easy adaptive adjustment. We use sparsification compression and incorporate EF. We denote $\mathbf{C}_{\delta}(\cdot)$ as the compression operator with the compression ratio $\delta$ and $\Delta_{t}^{i}$ as the compressed update at $t$-th iteration in $i$-th worker. $\mathbf{e}_{t}^{i}$ represents the error term of the $i$-th worker at the $t$-th iteration, storing the discrepancy between the compressed updates and the true ones. The update rules for D-SGD with gradient compression are $\Delta_{t}^{i}=\mathbf{C}_{\delta}(\mathbf{e}_{t}^{i}+\gamma\mathbf{g}_{t}^{i})$, $\mathbf{e}_{t+1}^{i}=\mathbf{e}_{t}^{i}+\gamma\mathbf{g}_{t}^{i}-\Delta_{t}^{i}$ and $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{1}{n}\sum_{i=1}^{n}\Delta_{t}^{i}$. As shown in Fig. 3, D-SGD with gradient compression retains a serial process, but reduces transmission time compared to D-SGD.
Figure 3: The running timelines for D-SGD and D-SGD with communication optimziation strategies. Both D-SGD and D-SGD with gradient compression are serial processes, with gradient compression reducing transmission time. D-SGD with delayed aggregation and DD-EF-SGD operate in parallel.
2.2.3 D-SGD with Delayed Aggregation
@@
Delayed aggregation DGA addresses communication delays by accounting for the delay staleness $\tau$ between the computation of gradients and their application to the model update stich2019error. $\mathbf{g}_{t-\tau}^{i}$ represents the gradient computed by worker $i$ at the $(t-\tau)$-th iteration and the update rule shows as $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}$. Unlike standard D-SGD, delayed aggregation enables pipelining: the system performs the $t$-th computation concurrently with the transmission of gradients from the $(t-\tau)$-th iteration (as shown in Fig. 3). While the physical transmission duration remains unchanged, it is effectively hidden behind the computation. When further combined with gradient compression (DD-EF-SGD), the transmission time is reduced.
@@ 2.3 Virtual Sequence
The Virtual Sequence method stich2018sparsified; mania2017perturbed is an analytical tool designed to simplify the convergence analysis of D-SGD variants like D-SGD with gradient compression. It defines an auxiliary sequence, $\tilde{\mathbf{x}}_{t}$, which evolves according to an idealized update rule and serves as a reference for the actual iterates, $\mathbf{x}_{t}$, accounting for noise introduced by compression and delay. We define the noise term as $B_{t}$, then we have $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$. We generalize updates as $\mathbf{v}_{t}$, (i.e., $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\mathbf{v}_{t}$) and $B_{t}$ can be computed in $B_{t+1}=B_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\mathbf{v}_{t}$. Then we have $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\textbf{g}_{t}^{i}$. While effective for simple variants, this tool fails for complex coupled systems like DD-EF-SGD. The failure arises because the noise term in DD-EF-SGD involves a recursive coupling of compression errors and delay staleness. This dependency makes the noise term mathematically intractable under standard noise assumptions, necessitating a new analytical method to decouple these factors.